OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (secret-ninja):

I think you already asked this one before....

OpenStudy (anonymous):

idk but i need help what isthe formula

OpenStudy (secret-ninja):

For slope: (y-y1) = m (x-x1) = m Standard slope form: y=mx+b

OpenStudy (secret-ninja):

x-intercept of 6 = (6, 0) y-intercept of 5 = (0, 5)

OpenStudy (anonymous):

ok thank you

OpenStudy (secret-ninja):

No problem. :)

OpenStudy (anonymous):

so m=6?

OpenStudy (secret-ninja):

Actually, its 6/5 which is 1.2 in decimal form

OpenStudy (secret-ninja):

If you subtract 0 it stays the same.

OpenStudy (anonymous):

A. y = 6x + 5 B. y = -5/6 x + 5 C. y = -6x + 5 D. y = 5/6 x + 5

OpenStudy (anonymous):

so its not a or c

OpenStudy (anonymous):

so d?

OpenStudy (anonymous):

confused now .-.

OpenStudy (anonymous):

is it d or not

OpenStudy (mathstudent55):

Here is the explanation from the beginning. The slope-intercept form of the equation of a line is: \(y = mx + b\), where m is the slope, and b is the y-coordinate of the y-intercept. You are given the intercepts, so using those two points, you can find m, the slope. \(m = \dfrac{y_2 - y_1}{x_2 - x_1} \)

OpenStudy (mathstudent55):

The x-intercept gives you point: (6, 0) The y-intercept gives you point: (0, 5) Now we find the slope: \(x = \dfrac{0 - 5}{6 - 0} = - \dfrac{5}{6}\)

OpenStudy (anonymous):

is it d? and ok its b?

OpenStudy (mathstudent55):

Now that we have the slope, and we know the y-intercept is 5, we just use the slope-intercept equation and plug in our values: \(y = mx + b\) \(y = -\dfrac{5}{6}x + 5\)

OpenStudy (mathstudent55):

BTW, two responses above, I meant to write m = the fraction of the slope, not x.

OpenStudy (anonymous):

i under stand but is b correct???